# references.bib

### v2

@@ Abstract.
A growing body of work has shown that AI-assisted methods — leveraging large language models (LLMs), social choice methods, and collective dialogues — can help reduce polarization and foster common ground in controlled lab settings. But what can these approaches contribute in real-world contexts? We present a case study applying these techniques to find common ground between Israeli and Palestinian peacebuilders in the period following October 7th, 2023. From April to July 2024 an iterative deliberative process combining LLMs, bridging-based ranking, and collective dialogues was conducted in partnership with the Alliance for Middle East Peace. More than 100 civil society peacebuilders participated including Israeli Jews, Palestinian citizens of Israel, and Palestinians from the West Bank and Gaza. The process culminated in a set of collective statements, including joint demands to world leaders, with at least 84% agreement from participants on each side. In this paper we review the mechanics and implementation of the process, discuss results and learnings, and highlight open problems that warrant future work.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ACM Conference on Fairness, Accountability, and Transparency; June 03–05, 2025; Athens, Greeceisbn: 978-1-4503-XXXX-X/18/06copyright: nonefootnotetext: Corresponding authors: andrew@remesh.org, bakker@mit.edu
Figure 1. The effort consisted of four dialogue cycles with peacebuilders: three uninational dialogues cycles conducted respectively with Israeli Jews, Palestinians from the West Bank and Gaza, and Palestinian citizens of Israel, followed by a final joint dialogue involving all three groups. Each dialogue cycle involved a process to find common ground among participants followed by a deliberation on the results of that process. The common ground process started with a collective dialogue, then bridging-based ranking was used to identify common ground ’bridging statements’ that were distilled into articulate ’collective statements’ via LLM and reviewed by human experts before being shared back with participants for a final vote. The method is described in detail in Section 3.
@@ 1. Introduction
This paper documents an effort to find common ground between Israeli and Palestinian peacebuilders amidst deeply entrenched divisions that were amplified in the immediate aftermath of October 7th, 2023 and shaped by the broader context of protracted asymmetric conflict. From April to July 2024, an iterative deliberative process that combined collective dialogues (ovadya2023generative), bridging-based ranking (ovadya2022bridging), and LLMs (goldberg2024ai) was carried out in partnership with the Alliance for Middle East Peace (ALLMEP). More than 100 civil society peacebuilders participated in the process including Israeli Jews, Palestinian citizens of Israel, and Palestinians from the West Bank and Gaza, representing diverse lived experiences and interests. The process produced a set of collective statements (Table 3) reflecting points of common ground — including five statements to world leaders, and five to residents of the region — with each statement supported by no fewer than 84% of peacebuilders on any side.
@@
In recent years, there have been several proposals using algorithmic methods to address polarization and find common ground. Bridging-based ranking, a class of recommendation algorithms that aims to surface content that builds trust and mutual understanding across divides (ovadya2022bridging; ovadya2023), was first implemented on Australian policy forum YourView in 2012 (van-gelder2012), and later gained prominence when independently implemented on Polis (small2021polis; polisMathCommit8394f1f) and the Community Notes feature on X (then Twitter) (wojcik2022bird). Large-scale, online collective dialogues (ovadya2023generative) that incorporate matrix completion methods to reduce the burden of preference elicitation have been used by the United Nations for peacebuilding since 2020 (bilich2019faster; masood2022using; irwin2021using). Most recently, there have been several proposals to use LLMs and social choice mechanisms to find common ground and make collective decisions on divisive issues (bakker2022fine; small2023opportunities; fish2023generative; tessler2024ai; de2025supernotes).
@@
Most study of these methods — particularly bridging-based ranking and LLM-based approaches to articulating common ground — has occurred in lab settings, offline, or for relatively low-stakes contexts. For example, the approach used in this case study builds on a specific integration of these methods that was initially developed to produce common ground policy as part of OpenAI’s Democratic Inputs to AI program (konya2023dem), and subsequently refined for use in a study on AI alignment (konya2024chain). In contrast, we present a case study using a combination of these methods to find common ground between cross-border stakeholders amid active armed violence. This required us to address a range of real-world challenges related to trust, legitimacy, language barriers, and a constantly-changing situation. We believe this work may be the first to examine the use of LLMs to bridge divides during a high-stakes real-world conflict.
@@
  • Impartiality: Drawing on best practices from peacebuilding (un_peacekeeping_2008) and mediation (gaffney2022impartiality), we aimed for both our process and technology to be unbiased and fair: eliciting and representing participant perspectives without bias, and enabling equitable participation and influence over outcomes.
@@
In this vein, we also emphasize that this paper constitutes a case study, not research. The project was not undertaken for academic purposes, but as a good-faith attempt to support real-world peace processes. Opportunities for controlled experimentation and evaluation to inform design decisions were limited. As such, our approach reflects a pragmatic reconciliation of established techniques with the contextual constraints of a complex sociopolitical situation. Nonetheless, we believe the project can inform future academic work on the design and use of technology in conflict settings and are thus, with the permission of ALLMEP, publishing this case study.
The paper is structured as follows. Section 2 reviews situational and technical context which motivated the project. Section 3 documents the method used to find common ground and explains why key design decisions were made. Section 4 reports the main insights and statements (representing common ground) that resulted from the process. Finally, in Section 5 we discuss ethical considerations, the limitations of the process, and directions we intend to prioritize in future work.
@@ 2.1. Situational
In the context of geopolitical conflicts, “track 1 diplomacy” involving formal negotiations between governments is often supplemented or supplanted by “backchannel” dialogue involving unofficial non-governmental representatives from each side. Such backchannel dialogues may involve government officials participating in an unofficial capacity (“track 1.5 diplomacy”) or may not involve governments at all (“track 2 diplomacy”) (montville1991track). Often, these dialogues provide a more open environment to build trust and discuss tough issues. Furthermore, they can enable cross-border communication when formal diplomatic channels are closed. Civil society peacebuilders typically play a significant role in convening and contributing to these dialogues.
In early 2024, following the events of October 7th, 2023, and within the broader context of protracted asymmetrical conflict, tensions between Israeli and Palestinian peacebuilders had understandably intensified. Shaped not only by the recent escalation in violence, but also by decades of structural disparity, open hostility, and deep-seated mistrust, the substrate to discuss tough issues had grown fragile [^1] We intentionally chose to use politically neutral language and not impose a characterization of the lived experiences of Palestinians and Israelis, including our own. This includes deliberately omitting certain terms, such as “occupation,” that have widely recognized legal and political significance under international law and are central to the lived experiences of many participants in this process. We believe it is consistent with the spirit of this work that any political positions conveyed by this otherwise-technical paper should not be ours, but those manifest in the points of common ground found among the peacebuilders who participated in this effort (Table 3).. In this context, we worked with the Alliance for Middle East Peace (ALLMEP) — a network of civil society peacebuilders in the region — with the goal of strengthening that substrate. This initiative sought to create a foundation for meaningful dialogue grounded in shared humanity and collective action. To this end, we sought to help peacebuilders find common ground. But the situation presented a few real-world challenges.
@@
  • Trust: Skepticism and mistrust, stemming from historical experiences of injustice, violence and prolonged conflict, posed significant barriers to engagement. This meant that before we could reasonably expect a willingness to participate in joint dialogue, we would need to elevate trust in good faith partnership. It also meant that the process itself would have to hold up to scrutiny that the “other side” did not have a mechanism to influence the results to their advantage.
  • Asymmetry: We expected that, by default, there would be an unequal number of representatives participating from each side, and that it would be not be in the spirit of inclusion to limit the number of participants from each side to achieve balance. It was crucial to design a process that accounted for structural asymmetries, ensuring that even with unequal representation, all “sides” held equitable influence on the outcomes.
@@ 2.2.1   Collective dialogues
An online collective dialogue process is an iterative back-and-forth exchange between a moderator and participants. Participants typically join by clicking a link, and fill out demographic information as they enter the dialogue. During each turn of the dialogue, participants are sent either a read-only message, a standard survey question, or an open-ended prompt that kicks off a collective response process. During a collective response process (ovadya2023generative), participants submit statements in response to the prompt, then evaluate statements submitted by others. The evaluation step serves two purposes: it exposes participants to others’ views to help them understand each other better, and it elicits data for quantifying each statement’s representativeness. On the collective dialogue platform Remesh, which we used for this project, two types of evaluations are elicited; agreement votes, and pair choice votes. Then, since participants can’t feasibly vote on all statements submitted to each prompt, a matrix completion method (konya2022elicitation; bilich2019faster) is used to convert a sparse vote sampling into a complete participant vote matrix. The percent agreement each participant segment has with each statement is then computed by combining the complete vote matrix with the demographic data of the participants.
During a live collective dialogue process, all participation is simultaneous and each collective response process takes a few minutes. When each collective response process completes, each participant sees the fraction of all participants agreeing with their response along with a representative subset of responses from the group. The moderator sees live results which they can use to decide if they need to pivot from the pre-programmed discussion guide to ask additional questions on an issue. After a live dialogue is complete it can be left open for asynchronous participation so those who could not make the live event can still participate. Collective dialogues can also be run asynchronously from the start if the logistics of a live dialogue are not feasible for the situation. In either case, a collective dialogue typically includes on the order of ten collective response prompts, and takes participants between 15-60 minutes.
@@
The United Nations has run collective dialogues on Remesh — often referred to as “Digital Dialogues” — as part of peacebuilding efforts since 2020 (futuringPeaceDigitalDialoguesYemen). These collective dialogues are sometimes conducted by special political missions in support of track 1 diplomatic processes, as was done by Stephanie Williams during ceasefire negotiations in Libya (UNLibya; irwin2021using). In other cases, they serve less formal track 1.5 or track 2 activities associated with local field presences. Overall, the UN has now run more than 60 collective dialogues across the world including in Yemen (futuringPeaceDigitalDialoguesYemen) , Iraq (UNIraq) , Lebanon (UNLebanon), Haiti (UNHati), Bolivia (UNBolivia; UNBolivia2), and Bahrain (UNBahrain).
2.2.2   Bridging-based ranking
@@
The term bridging-based ranking refers to a class of ranking or recommendation algorithms that aim to “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation” (ovadya2023). Most commonly, this qualitative goal is operationalized as “diverse approval,” meaning that items are ranked by the degree to which they are approved of or valued by clusters of participants who would normally be expected to disagree with each other (in other words, ranking by diverse approval surfaces a form of “common ground”). Implementations of this core intuition differ, for example with respect to whether clusters are stipulated or learned from the data (small2021polis), and how diversity is quantified. Below, we review two existing approaches which we used in this project.
Let $P$ denote a set of people, $\mathcal{G}$ denote a set of groups (that is, a set of subsets of $P$), and $S$ denote a set of items to be ranked, which in our context are natural language statements.
@@
  • Perhaps the simplest approach is max-min agreement, where items are ranked by the minimum rate of agreement they receive across each group. Formally, for a group $G\in\mathcal{G}$ and statement $s\in S$, let $a(s,G)$ denote the proportion of people in group $G$ that agree with statement $s$. Under the max-min agreement approach, statements in $s\in\mathcal{S}$ are ranked according to the metric
@@
    This metric has been used previously in collective dialogue processes run on Remesh (konya2023dem; konya2024chain).
  • Another prominent approach to bridging is the algorithm used by the Community Notes feature on the social media platform X, which elicits crowd-sourced notes that add context to posts, but only shows them publicly if (to simplify slightly) they are rated as helpful by people on both sides of the political spectrum. To achieve this, the algorithm fits a latent factor model to the observed agreement votes, which provides for each note a factor corresponding to its slant along the most salient dimension of disagreement (wojcik2022bird). Formally, let $v(s,p)\in\{1,-1\}$ denote whether person $p$ said they respectively agreed or disagreed with statement $s$. The Community Notes algorithm then fits the following model
@@
    The global intercept $\mu$ captures the overall likelihood of any voter to agree [^2] In the Community Notes implementation used for posts on X, the primary vote reflected in $\hat{v}(s,p)$ is not “agree”, but “helpful”. While the math is the same, here we describe the algorithm in terms of “agree” votes in order to maintain consistency across descriptions of the methods used in this paper. with any statement, the user intercept $i_{p}$ captures voter $p$’s overall likelihood to agree with any statement, and the note intercept $i_{s}$ captures the voting population’s overall likelihood to agree with statement $s$. The last term $\mathbf{f}_{p}\cdot\mathbf{f}_{s}$ represents the product of the voter and statement factors, [^3] These factors can be scalars that capture only the primary axis of voter variance, as in the original Birdwatch paper (wojcik2022bird), or vectors that capture multi-dimensional factors of vote variance., respectively. Model parameters are fit by minimizing a regularized least-squared loss via gradient descent over the dataset of observed votes. Once learned, the note and user factors map statements and voters along learned axes spanning “sides” of division, while the note intercept $i_{s}$ gives a side-independent measure of agreement. Because the note intercept $i_{s}$ is side-independent, we can use it to find statements that bridge across people on different sides.
2.2.3   Language models
Large language models (LLMs) are increasingly being recognized for their potential to address societal challenges by enhancing collaboration and fostering collective intelligence. These models offer new ways to process and synthesize diverse viewpoints, enabling the articulation of shared perspectives even in deeply divided groups (argyle2023leveraging; fish2023generative; tessler2024ai). For example, in collective dialogue systems, LLMs can distill participant responses into coherent statements that reflect areas of agreement.
Beyond synthesizing viewpoints and identifying areas of agreement, LLMs can play a critical role in overcoming language barriers. They offer real-time translation, text-to-speech and summarization capabilities, ensuring participants can engage fully in their native languages. This functionality helps make deliberative processes more inclusive, equitable, and accessible (burton2024large). However, careful implementation is essential to avoid risks such as bias and misrepresentation.

### v3

# Using Collective Dialogues and AI to Find Common Ground Between Israeli and Palestinian Peacebuilders

@@ Abstract.
A growing body of work has shown that AI-assisted methods — leveraging large language models, social choice methods, and collective dialogues — can help navigate polarization and surface common ground in controlled lab settings. But what can these approaches contribute in real-world contexts? We present a case study applying these techniques to find common ground between Israeli and Palestinian peacebuilders in the period following October 7th, 2023. From April to July 2024 an iterative deliberative process combining LLMs, bridging-based ranking, and collective dialogues was conducted in partnership with the Alliance for Middle East Peace. Around 138 civil society peacebuilders participated including Israeli Jews, Palestinian citizens of Israel, and Palestinians from the West Bank and Gaza. The process resulted in a set of collective statements, including demands to world leaders, with at least 84% agreement from participants on each side. In this paper, we document the process, results, challenges, and important open questions.
Peace, Conflict, Collective Response System, Collective Dialogue, Bridging, Bridging-Based Ranking
journalyear: 2025conference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732022isbn: 979-8-4007-1482-5/2025/06ccs: Human-centered computing Computer supported cooperative workccs: Applied computing Voting / election technologies
@@ 1. Introduction
This paper documents an effort to find common ground between Israeli and Palestinian peacebuilders amidst deeply entrenched divisions that were amplified in the immediate aftermath of October 7th, 2023 and shaped by the broader context of protracted asymmetric conflict. From April to July 2024, an iterative deliberative process that combined collective dialogues (Ovadya, 2023), bridging-based ranking (Ovadya, 2022), and large language models (LLMs) (Goldberg et al., 2024) was carried out in partnership with the Alliance for Middle East Peace (ALLMEP). More than 100 civil society peacebuilders participated in the process including Israeli Jews, Palestinian citizens of Israel, and Palestinians from the West Bank and Gaza, representing diverse lived experiences and interests. The process produced a set of collective statements (Table 3) reflecting points of common ground — including five statements to world leaders, and five to residents of the region — with each statement supported by at least 84% of peacebuilders on each side.
@@
In recent years, there have been several proposals using algorithmic methods to address polarization and find common ground. Bridging-based ranking, a class of recommendation algorithms that aims to surface content that builds trust and mutual understanding across divides (Ovadya, 2022; Ovadya and Thorburn, 2023), was first implemented on Australian policy forum YourView in 2012 (van Gelder, 2012), and later gained prominence when independently implemented on Polis (Small et al., 2021; Small, 2015) and the Community Notes feature on X (then Twitter) (Wojcik et al., 2022). Large-scale, online collective dialogues (Ovadya, 2023) that incorporate matrix completion methods to reduce the burden of preference elicitation have been used by the United Nations for peacebuilding since 2020 (Bilich et al., 2023; Masood Alavi et al., 2022; Alavi et al., 2022). Most recently, there have been several proposals to use LLMs and social choice mechanisms to find common ground and make collective decisions on divisive issues (Bakker et al., 2022; Small et al., 2023; Fish et al., 2025; Tessler et al., 2024; De et al., 2025).
@@
Most study of these methods — particularly bridging-based ranking and LLM-based approaches to articulating common ground — has occurred in lab settings, offline, or for relatively low-stakes contexts. For example, the approach used in this case study builds on a specific integration of these methods that was initially developed to produce common ground policy as part of OpenAI’s Democratic Inputs to AI program (Konya et al., 2023), and subsequently refined for use in a study on AI alignment (Konya et al., 2024). In contrast, we present a case study using a combination of these methods to find common ground between cross-border stakeholders amid active armed violence. This required us to address a range of real-world challenges related to trust, legitimacy, language barriers, and a constantly-changing situation. We believe this work may be the first to examine the use of LLMs to bridge divides during a high-stakes real-world conflict.
@@
  • Impartiality: Drawing on best practices from peacebuilding (United Nations, 2008) and mediation (Gaffney, 2022), we aimed for both our process and technology to be unbiased and fair: eliciting and representing participant perspectives without bias, and enabling equitable participation and influence over outcomes.
@@
In this vein, we emphasize this paper presents a case study from a project aimed at supporting real-world peace processes, not academic research, which limited opportunities for controlled experimentation or formal evaulation. As such, our approach reflects a pragmatic reconciliation of established techniques with the constraints of a complex sociopolitical situation. Nonetheless, we believe the project can inform future academic work on the design and use of technology in conflict settings and are thus, with the permission of ALLMEP, publishing this case study.
The paper is structured as follows. Section 2 reviews situational and technical context which motivated the project. Section 3 documents the method used to find common ground and explains key design decisions. Section 4 reports the main insights and statements (representing common ground) that resulted from the process. Finally, in Section 5 (and related appendices) we discuss ethical considerations, the limitations of the process, and directions we intend to prioritize in future work.
Figure 1. The effort consisted of four dialogue cycles with peacebuilders: three uninational dialogues cycles conducted respectively with Israeli Jews, Palestinians from the West Bank and Gaza, and Palestinian citizens of Israel, followed by a final joint dialogue involving all three groups. Each dialogue cycle involved a process to find common ground among participants followed by a deliberation on the results of that process. The common ground process started with a collective dialogue, then bridging-based ranking was used to identify common ground “bridging statements” that were distilled into articulate “collective statements” via LLM and reviewed by human experts before being shared back with participants for a final vote. The method is described in detail in Section 3.
@@ 2.1. Situational
In the context of geopolitical conflicts, “track 1 diplomacy” involving formal negotiations between governments is often supplemented or supplanted by “backchannel” dialogue involving unofficial non-governmental representatives from each side. Such backchannel dialogues may involve government officials participating in an unofficial capacity (“track 1.5 diplomacy”) or may not involve governments at all (“track 2 diplomacy”) (Montville, 1991). Often, these dialogues provide a more open environment to build trust and discuss tough issues. Furthermore, they can enable cross-border communication when formal diplomatic channels are closed. Civil society peacebuilders typically play a significant role in convening and contributing to these dialogues.
In early 2024, following the events of October 7th, 2023, and within the broader context of protracted asymmetrical conflict, tensions between Israeli and Palestinian peacebuilders had understandably intensified. Shaped not only by the recent escalation in violence, but also by decades of structural disparity, open hostility, and deep-seated mistrust, the substrate to discuss tough issues had grown fragile [^1] In this paper, we intentionally chose to use politically neutral language and not impose a characterization of the lived experiences of Palestinians and Israelis, including our own. This includes deliberately omitting certain terms, such as “occupation,” that have widely recognized legal and political significance under international law and are central to the lived experiences of many participants in this process. We believe it is consistent with the spirit of this work that any political positions conveyed by this otherwise-technical paper should not be ours, but those manifest in the points of common ground found among the peacebuilders who participated in this effort (Table 3).. In this context, we worked with the Alliance for Middle East Peace (ALLMEP) — a network of civil society peacebuilders in the region — with the goal of strengthening that substrate. This initiative sought to create a foundation for meaningful dialogue grounded in shared humanity and collective action. To this end, we sought to help peacebuilders find common ground. But the situation presented a few real-world challenges.
@@
Trust: Skepticism and mistrust, stemming from historical experiences of injustice, violence and prolonged conflict, posed significant barriers to engagement. This meant that before we could reasonably expect willingness to participate in joint dialogue, we would need to elevate trust in good faith partnership. It also meant that the process itself needed to hold up to scrutiny that the “other side” did not have a mechanism to influence the results to their advantage.
Asymmetry: We expected, by default, there would be an unequal number of representatives participating from each side, yet it would be anti-inclusionary to turn away participants from a side to achieve balance. It was crucial to design a process that accounted for structural asymmetries, ensuring that even with unequal representation, all “sides” held equitable influence on the outcomes.
@@ 2.2.1. Collective dialogues
An online collective dialogue process is an iterative back-and-forth exchange between a moderator and participants. Participants join by clicking a link, and fill out demographic information as they enter the dialogue. During each turn of the dialogue, participants are sent either a read-only message, a standard survey question, or an open-ended prompt that kicks off a collective response process. During a collective response process (Ovadya, 2023), participants submit statements in response to the prompt, then evaluate statements submitted by others. The evaluation step serves two purposes: it exposes participants to others’ views to help them better understand each other, and it elicits data for quantifying each statement’s representativeness. On the collective dialogue platform used for this project (Remesh), two types of evaluations are elicited; agreement votes, and pair choice votes. To maximize the number of dialogue turns viable within a fixed budget of participant time, and decouple votes per participant needed from the number of statements submitted, votes are sparsely sampled and a matrix completion method (Konya et al., 2022; Bilich et al., 2023) is used to convert a sparse vote sampling into a complete participant vote matrix. The percent agreement each participant segment has with each statement is then computed by combining the complete vote matrix with the demographic data of the participants. Collective dialogues can be run both live and asynchronously (see Appendix B.1). They typically include on the order of ten collective response prompts, and take participants between 15-60 minutes to complete.
@@
The United Nations has run collective dialogues on Remesh — often referred to as “Digital Dialogues” — as part of peacebuilding efforts since 2020 (United Nations DPPA Innovation, 2022g). These collective dialogues are sometimes conducted by special political missions in support of track 1 diplomatic processes, as was done by Stephanie Williams during ceasefire negotiations in Libya (United Nations DPPA Innovation, 2022f; Alavi et al., 2022). In other cases, they serve less formal track 1.5 or track 2 activities associated with local field presences. Overall, the UN has now run more than 60 collective dialogues across the world including in Yemen (United Nations DPPA Innovation, 2022g), Iraq (United Nations DPPA Innovation, 2022d), Lebanon (United Nations DPPA Innovation, 2022e), Haiti (United Nations DPPA Innovation, 2022c), Bolivia (United Nations DPPA Innovation, 2022b, h), and Bahrain (United Nations DPPA Innovation, 2022a).
2.2.2. Bridging-based ranking
@@
The term bridging-based ranking refers to a class of ranking or recommendation algorithms that aim to “increase mutual understanding and trust across divides, creating space for productive conflict, deliberation, or cooperation”‘(Ovadya and Thorburn, 2023). Most commonly, this qualitative goal is operationalized as “diverse approval,” meaning that items are ranked by the degree to which they are approved of or valued by clusters of participants who would normally be expected to disagree with each other (in other words, ranking by diverse approval surfaces a form of “common ground”). Implementations of this core intuition differ, for example with respect to whether clusters are stipulated or learned from the data (Small et al., 2021), and how diversity is quantified. Below, we review two existing approaches which we used in this project.
@@
Let $P$ denote a set of people, $\mathcal{G}$ denote a set of groups (that is, a set of subsets of $P$), and $S$ denote a set of items to be ranked, which in our context are natural language statements. Perhaps the simplest approach is max-min agreement, where items are ranked by the minimum rate of agreement they receive across each group. Formally, for a group $G\in\mathcal{G}$ and statement $s\in S$, let $a(s,G)$ denote the proportion of people in group $G$ that agree with statement $s$. Under the max-min agreement approach, statements in $s\in\mathcal{S}$ are ranked according to the metric
@@
This metric has been used previously in collective dialogue processes run on Remesh (Konya et al., 2023, 2024).
Another prominent approach to bridging is the algorithm used by the Community Notes feature on the social media platform X. It elicits crowd-sourced notes that add context to posts, but only shows them publicly if (to simplify slightly) they are rated as helpful by people on both sides of a learned political spectrum. To achieve this, the algorithm fits a latent factor model to the observed agreement votes, which provides for each note a factor corresponding to its slant along the most salient dimension of disagreement (Wojcik et al., 2022). Formally, let $v(s,p)\in\{1,-1\}$ denote whether person $p$ said they respectively agreed or disagreed with statement $s$. The Community Notes algorithm then fits the following model
@@
The global intercept $\mu$ captures the overall likelihood of any voter to agree [^2] In the Community Notes implementation used for posts on X, the primary vote reflected in $\hat{v}(s,p)$ is not “agree”, but “helpful”. While the math is the same, here we describe the algorithm in terms of “agree” votes in order to maintain consistency across descriptions of the methods used in this paper. with any statement, the user intercept $i_{p}$ captures voter $p$’s overall likelihood to agree with any statement, and the note intercept $i_{s}$ captures the voting population’s overall likelihood to agree with statement $s$. The last term $\mathbf{f}_{p}\cdot\mathbf{f}_{s}$ represents the product of the voter and statement factors, [^3] These factors can be scalars that capture only the primary axis of voter variance, as in the original Birdwatch paper (Wojcik et al., 2022), or vectors that capture multi-dimensional factors of vote variance., respectively. Model parameters are fit by minimizing a regularized least-squared loss via gradient descent over the dataset of observed votes. Once learned, the note and user factors map statements and voters along learned axes spanning “sides” of division, while the note intercept $i_{s}$ gives a side-independent measure of agreement. Because the note intercept $i_{s}$ is side-independent, we can use it to find statements that bridge across people on different sides.
2.2.3. Language models
Large language models (LLMs) are increasingly being recognized for their potential to address societal challenges by enhancing collaboration and fostering collective intelligence. These models offer new ways to process and synthesize diverse viewpoints, enabling the articulation of shared perspectives even in deeply divided groups (Argyle et al., 2023; Fish et al., 2025; Tessler et al., 2024). For example, in collective dialogue systems, LLMs can distill participant responses into coherent statements that reflect areas of agreement. Beyond synthesizing viewpoints and identifying areas of agreement, LLMs can play a critical role in overcoming language barriers. They offer real-time translation, text-to-speech and summarization capabilities, ensuring participants can engage fully in their native languages. This functionality helps make deliberative processes more inclusive, equitable, and accessible (Burton et al., 2024). However, careful implementation is essential to avoid risks such as bias and misrepresentation.